ORDER COMPLEXES OF NONCOMPLEMENTED LATTICES ARE NONEVASIVE

DMITRY KOZLOV

Abstract. We reprove and generalize in a combinatorial way the result of A. Björner [J. Comb. Th. A 30, 1981, pp. 90–100, Theorem 3.3], that order complexes of noncomplemented lattices are contractible, namely by showing that these simplicial complexes are in fact nonevasive, in particular collapsible.